two equal positive charges q1 = q2 = 2 microcoulomb are located at x = 0, y = 0.30 metre and x = 0, y = - 0.30 metre respectively. what are the magnitude and direction of the electric force q1 and q2 exert on a third charge Q4 microcoulomb at x = 0.40 metre, y = 0?


1
Expert's answer
2021-02-24T12:47:28-0500

"F_1=F_2=\frac{kq_1q_3}{r^2},"

"F_1=F_2=\frac{9\cdot 10^9\cdot 2\cdot 10^{-6}\cdot 4\cdot10^{-6}}{{0.3^2+0.4^4}}=0.288~\text{N},"

"F=2F_1cos( atan \frac 34)=\frac 85 F_1=0.46~ \text {N},"

the direction is the same as the direction of the positive x-axis.